If TC=0.01x2+5x+100 is the cost function for x items of a product, at what level of production, x is there minimum average cost? What is the minimum average cost?


1
Expert's answer
2021-11-03T12:35:29-0400

"T_c(x)=0.01x^2+5x+100\\T'(x)=0.02x+5\\"

At optimal point: "T'_c(x_0)=0"

"0.02x_0+5=0\\0.02x_0=-5\\x_0=-\frac{5}{0.02}"

"x_0=-25<0"

"x_0=0"

"T_c(x_0)=0+0+0+100"

therefore, the minimum average cost is 100